among the reading club members there are 8 more girls than boys.If there are 50 members in all ,how many are boys?

Answer by lwsshak3(11628)

among the reading club members there are 8 more girls than boys.If there are 50 members in all ,how many are boys?
***
let x=number of girls
50-x=number of boys
50-x=x+8
2x=42
x=21
50-x=29
number of boys=29